OpenStudy (vortish):

find the vertex of the parabola f(x)=x^2-4x-2

OpenStudy (anonymous):

complete the squares of RHS

OpenStudy (anonymous):

this is the formula: x = -b / (2a) if you plug x into your formula, you'll find the y coordinate

OpenStudy (anonymous):

y= f(x)=x^2-4x-2 y=x^2-4x-2 y=f(x)=x^2-4x+4 -6 y=(x-2)^2 -6 y+6=(x-2)^2 vertex is 2,-6

OpenStudy (vortish):

never done these kind of equations and im lost when you you put up formulas i learn better by seeing a problem worked out but i do like to try to help in the formation of the answer

OpenStudy (anonymous):

ok, you're -b = 4. 2a = 2. 4/2 = 2, so x = 2. Now plug that 2 back into the function you just posted.

OpenStudy (vortish):

so would f(x)=-6

OpenStudy (vortish):

it wants the answer as a ordered pair so would the pair be (2,-6)?

OpenStudy (anonymous):

yes, that is it. exactly :)

OpenStudy (anonymous):

its just a simple formula, do many problems.

OpenStudy (vortish):

ty odd for walking threw that with me im a right brain person i have to see it a few times to understand it but i also dont just want the answers handed to me i like to try to learn by doing some of the work to